Prepare 2L of Solution A given the following stock solutions.

Stock Solutions:

5 M NaCl

10% Ethanol

Solution A

250 mM NaCL

0.025% Ethanol

Your answer must include the volumes of 5M NaCl, 10% Ethanol, and water.
Answer #1

For calculating the volume of NaCl:

M1*V1 = M2*V2

5*V1 = 0.25*2

V1= 0.1 L

The volume of NaCl solution needed is 0.1 L

For calculating the volume of Ethanol:

C1*V1=C2*V2

10*V1=0.025*2

V1= 0.005 L

The volume of Ethanol needed is 0.005 L

For making the 2 L solution A, the volume of water needed is = 2 - (0.1+0.005) L

= 1.895 L